OpenStudy (anonymous):

What is the d value of the following arithmetic sequence? 5, 2, -1, -4, -7

OpenStudy (asnaseer):

look at what you need to do to each term in order to get to the next one.

OpenStudy (asnaseer):

so, for example, what do you need to do to 5 in order to get to 2?

OpenStudy (anonymous):

@asnaseer -3???

OpenStudy (asnaseer):

correct. now just double check that this works for all subsequent terms

OpenStudy (anonymous):

@asnaseer how do i check?

OpenStudy (asnaseer):

well the next term is 2. so if you subtract 3 from 2 what do you get?

OpenStudy (asnaseer):

1st term = 5 2nd term = 5 - 3 = 2 3rd term = 2 - 3 = ?

OpenStudy (anonymous):

@asnaseer i got it thanks!!! just to make sure (Find the six arithmetic means between 1 and 29) i got 5, 9, 13, 17, 21, 25

OpenStudy (asnaseer):

hang on - thats a different question you just listed? have you understood this question?

OpenStudy (anonymous):

yea

OpenStudy (asnaseer):

so what do you think 'd' represents in this question?

OpenStudy (anonymous):

idk

OpenStudy (asnaseer):

:/

OpenStudy (anonymous):

i know :(

OpenStudy (asnaseer):

it represents the common difference of an arithmetic sequence.

OpenStudy (anonymous):

ok.. but for the one i listed above you have to do something different to get the answer?

OpenStudy (asnaseer):

first of all I would recommend you look at this site to fully understand what an arithmetic sequence is: http://www.mathsisfun.com/algebra/sequences-sums-arithmetic.html for your new question - could you please post it as a new question in the list on the left